OpenStudy (anonymous):

What is the sum and how did you get it (-5) +(-8) +(-11) +... +-95. Possable answers a -1550 b -3100 c -1472 or d 300.

OpenStudy (anonymous):

Simplify the following: -5-8-11-95 Group the negative terms in -5-8-11-95 together and factor out the minus sign. -5-8-11-95 = -(5+8+11+95): -(5+8+11+95) Evaluate 5+8+11+95 using long addition. | 1 | | 9 | 5 | 1 | 1 | | 8 + | | 5 1 | 1 | 9: Answer: | | -119

OpenStudy (austinl):

\(\large{\sum\limits_{i=-5}^\infty a_{n-1}+(a_{n-1}-3)}\) I believe this is correct.

OpenStudy (anonymous):

That's what I kept getting but the possable answers are -1550,-3100, -1472, or 300